OpenStudy (iwanttogotostanford):

Given the function f(x) = x^2 and k = –1, which of the following represents a function opening downward? f(x) + k kf(x) f(x + k) f(kx)

OpenStudy (iwanttogotostanford):

@sleepyhead314 @wolf1728

OpenStudy (sleepyhead314):

the coefficient of x would have to be negative in order for the function to open downward

OpenStudy (wolf1728):

Figured I'd come here but sleepyhead seems to be ding quite well

OpenStudy (sleepyhead314):

since "k" is a negative number I would assume that they would like for us to multiply "x" with "k" somehow

OpenStudy (iwanttogotostanford):

i was thinking it'd be B or D?

OpenStudy (sleepyhead314):

D means that you would replace "kx" into the equation for "x" that would get you f(kx) = (-1x)^2 ^^ since this is squared, it would come out as a Positive so that cannot be right

OpenStudy (iwanttogotostanford):

i don't know what it would be...

OpenStudy (iwanttogotostanford):

@sleepyhead314

OpenStudy (sleepyhead314):

since it cannot be D then your other guess may be correct ;)

OpenStudy (wolf1728):

would it be k(fx)

OpenStudy (sleepyhead314):

yes ^

OpenStudy (sleepyhead314):

although the formatting is a bit confusing .-.

OpenStudy (wolf1728):

k = -1 f(x) = x^2 then k(fx) = -x^2

OpenStudy (iwanttogotostanford):

so it would be B??:-)

OpenStudy (wolf1728):

I think it is B

OpenStudy (sleepyhead314):

mhmmm

OpenStudy (wolf1728):

yes as sleepyhead said the formatting is weird

OpenStudy (iwanttogotostanford):

thank you both! I will have one more question up and then done!

OpenStudy (wolf1728):

Okay, I'll stay here